# All loop soft photon theorems and higher spin currents on the celestial sphere

### v1

@@ II All loop soft theorems in momentum space
In the presence of additional long-range interactions mediated by other massless fields, the soft factor coefficients $\mathbb{S}_{\text{em}}^{(n)}$ receive further universal corrections for all $n\geq 0$. For instance, the correction to $\mathbb{S}_{\text{em}}^{(0)}$ due to gravitational interactions was analyzed in Sahoo:2018lxl . In the context of the results presented above, even if we restrict attention to theories in which the only massless fields are the photon and the charged fields appearing as external hard particles, loop diagrams involving the same massless charged particles—specifically ladder and crossed-ladder topologies—can still generate additional contributions to $\mathbb{S}_{\text{em}}^{(n)}$ when three-point interaction vertices among the massless charged particles are allowed. This effect is discussed in a footnote of Appendix A for the case of scalar charged particle scattering. Consequently, the universality of the soft factors $\mathbb{S}_{\text{em}}^{(n)}$ for $n\geq 0$ is more restrictive than that of the leading soft factor $\mathbb{S}_{\text{em}}^{(-1)}$, as discussed in Krishna:2023fxg .
@@ III Soft theorems in conformal primary basis
for $j=1/2,1,3/2,\cdots$. In the above equations we have kept the helicity and all other internal quantum numbers implicit because the soft factors depend only on the electric charges and the momenta of the external hard particles. The soft operator $\mathcal{S}^{\Delta}(w,\bar{w})$ for positive helicity outgoing soft photon has been defined via the Mellin transformation of the photon annihilation operator $a(k,\sigma=+1)$ along with the conformal soft limit Donnay:2018neh taken inside the celestial correlation functions as [^7] Here we used the following result of Mellin transformation $\int_{0}^{\infty}d\omega\,\omega^{s-1}\omega^{n}(\ln\omega)^{k}=\frac{(-1)^{k}k!}{(s+n)^{k+1}}$, as derived in Krishna:2023ukw . One may be tempted to resum all the multi-order poles in the scaling dimensions and interpret the resummed Mellin-space Ward identity as a simple pole in $\Delta$ with anomalous dimensions. However, the resulting structure of the sum appears to be of the form $\sum_{n=0}^{\infty}(\Delta+n)^{n+2}$ in Mellin space, or equivalently exhibits the integral behavior $\int_{0}^{\infty}d\omega\omega^{\omega+\Delta-2}$, which falls into the category of the “Sophomore’s Dream”, for which no closed analytical result is known.
@@ IV $\mathcal{O}(\ln\omega)$ soft theorem as Ward identity and the dipole current operator
$\left\langle\mathcal{S}^{0}(w,\bar{w})\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle=\sum_{a=1}^{N}\frac{\eta_{a}{e_{a}}^{2}}{w-z_{a}}\Bigg[\sum_{\begin{subarray}{c}b=1\\ b\neq a\end{subarray}}^{N}e_{b}\,\frac{\bar{w}-\bar{z}_{b}}{\bar{z}_{a}-\bar{z}_{b}}\Bigg]\\ \left\langle\phi^{(\eta_{a})}_{\Delta_{a}-1}(z_{a},\bar{z}_{a},e_{a})\prod_{\begin{subarray}{c}i=1\\ i\neq a\end{subarray}}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle$
@@ V Conformal transformation law of the dipole operator
Under the $SL(2,\mathbb{R})_{\text{R}}$ transformation $\bar{z}\rightarrow\bar{z}^{\prime}=\frac{a\bar{z}+b}{c\bar{z}+d}$ with $a,b,c,d\in\mathbb{R}$ and adbc=1ad-bc=1, the invariance of correlation function (21) requires the dipole current $\bar{d}(\bar{w},\bar{z})$ to transform in the following way,
@@
Now specializing to the infinitesimal translation, dilatation and special conformal transformation (SCT)888 $\displaystyle\text{Translation:}\quad\begin{pmatrix}1&b\\ 0&1\end{pmatrix},\qquad\text{Dilatation:}\quad\begin{pmatrix}\lambda^{\frac{1}{2}}&0\\ 0&\lambda^{-\frac{1}{2}}\end{pmatrix},\qquad\text{SCT:}\quad\begin{pmatrix}1&0\\ -c&1\end{pmatrix}$, (30) we get the following commutation relations
@@
where $j=\frac{1}{2}$, $m=\pm\frac{1}{2}$ and the generators of translation, dilatation and SCT are denoted by $\bar{L}_{-1},\,\bar{L}_{0}$ and $\bar{L}_{1}$ respectively.
The transformation law (29) shows that the doublet $\left(\bar{d}^{\frac{1}{2}}_{\frac{1}{2}}(\bar{z}),\bar{d}^{\frac{1}{2}}_{-\frac{1}{2}}(\bar{z})\right)$ transform in the tensor product of two $SL(2,\mathbb{R})_{\text{R}}$ representations one of which is a conformal highest weight representation with weight $\bar{h}=1/2$ and the other one is a two dimensional spin-$\frac{1}{2}$ representation. Under the spin-$\frac{1}{2}$ representation the component fields $\bar{d}^{\frac{1}{2}}_{\pm\frac{1}{2}}(\bar{z})$ have weights $\pm\frac{1}{2}$. To see this at the level of infinitesimal transformation law, we denote the generator in the highest weight ($\bar{h}=\frac{1}{2}$) representation by $\bar{L}_{n}^{\bar{h}}$ and the generator in the spin-$\frac{1}{2}$ representation by $\bar{L}_{n}^{\bar{\sigma}}$. Now one can easily verify that the following commutation relations
@@ VII Correlation functions of dipole currents and matter fields
The dipole operator $\bar{d}^{\frac{1}{2}}_{\alpha}$ has scaling dimension $1/2$. Therefore conformal invariance determines the two point function to be  [^10] In terms of the original field $\bar{d}(\bar{w},\bar{z})$ this is nothing but $\langle d(\bar{w}_{1},\bar{z}_{1})\bar{d}(\bar{w}_{2},\bar{z}_{2})\rangle=k\frac{\bar{w}_{1}-\bar{w}_{2}}{\bar{z}_{1}-\bar{z}_{2}}$. This form of the two point function is uniquely determined by the $SL(2,\mathbb{R})_{\text{R}}$ transformation law (27).
@@
where $k$ is a real number which can be zero also. However, as we will see, we can take $k$ to be nonzero without changing the single soft photon theorems. Our current understanding of the role of dipole symmetry in the loop level soft theorems does not allow us to determine the value of $k$. This is discussed in more detail in section-XI.

### v2

@@ II All loop soft theorems in momentum space
In the presence of additional long-range interactions mediated by other massless fields, the soft factor coefficients $\mathbb{S}_{\text{em}}^{(n)}$ receive further universal corrections for all $n\geq 0$. For instance, the correction to $\mathbb{S}_{\text{em}}^{(0)}$ due to gravitational interactions was analyzed in Sahoo:2018lxl . In the context of the results presented above, even if we restrict attention to theories in which the only massless fields are the photon and the charged fields appearing as external hard particles, loop diagrams involving the same massless charged particles, specifically ladder and crossed-ladder topologies, can still generate additional contributions to $\mathbb{S}_{\text{em}}^{(n)}$ when three-point interaction vertices among the massless charged particles are allowed. This effect is discussed in a footnote of Appendix A for the case of scalar charged particle scattering. Consequently, the universality of the soft factors $\mathbb{S}_{\text{em}}^{(n)}$ for $n\geq 0$ is more restrictive than that of the leading soft factor $\mathbb{S}_{\text{em}}^{(-1)}$, as discussed in Krishna:2023fxg .
@@ III Soft theorems in conformal primary basis
for $j=1/2,1,3/2,\cdots$. In the above equations we have kept the helicity and all other internal quantum numbers of hard particles implicit because the soft factors depend only on the electric charges and the momenta of them. The soft operator $\mathcal{S}^{\Delta}(w,\bar{w})$ for positive helicity outgoing soft photon has been defined via the Mellin transformation of the photon annihilation operator $a(k,\sigma=+1)$ along with the conformal soft limit Donnay:2018neh taken inside the celestial correlation functions as [^7] Here we used the following result of Mellin transformation $\int_{0}^{\infty}d\omega\,\omega^{s-1}\omega^{n}(\ln\omega)^{k}=\frac{(-1)^{k}k!}{(s+n)^{k+1}}$, as derived in Krishna:2023ukw . One may be tempted to resum all the multi-order poles in the scaling dimensions and interpret the resummed Mellin-space Ward identity as a simple pole in $\Delta$ with anomalous dimensions. However, the resulting structure of the sum appears to be of the form $\sum_{n=0}^{\infty}(\Delta+n)^{n+2}$ in Mellin space, or equivalently exhibits the integral behavior $\int_{0}^{\infty}d\omega\omega^{\omega+\Delta-2}$, which falls into the category of the “Sophomore’s Dream”, for which no closed analytical result is known.
@@ IV $\mathcal{O}(\ln\omega)$ soft theorem as Ward identity and the dipole current operator
$\left\langle\mathcal{S}^{0}(w,\bar{w})\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle=\sum_{a=1}^{N}\frac{\eta_{a}{e_{a}}^{2}}{w-z_{a}}\Bigg[\sum_{\begin{subarray}{c}b=1\\ b\neq a\end{subarray}}^{N}e_{b}\,\frac{\bar{w}-\bar{z}_{b}}{\bar{z}_{a}-\bar{z}_{b}}\Bigg]\left\langle\phi^{(\eta_{a})}_{\Delta_{a}-1}(z_{a},\bar{z}_{a},e_{a})\prod_{\begin{subarray}{c}i=1\\ i\neq a\end{subarray}}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle$ (20)
@@ V Conformal transformation law of the dipole operator
Under the $SL(2,\mathbb{R})_{\text{R}}$ transformation $\bar{z}\rightarrow\bar{z}^{\prime}=\frac{a\bar{z}+b}{c\bar{z}+d}$ with $a,b,c,d\in\mathbb{R}$ and adbc=1ad-bc=1, the invariance of the correlation function (22) requires the dipole current $\bar{d}(\bar{w},\bar{z})$ to transform in the following way,
@@
and for infinitesimal translation, dilatation and special conformal transformations (SCT)888 $\displaystyle\text{Translation:}\quad\begin{pmatrix}1&b\\ 0&1\end{pmatrix},\qquad\text{Dilatation:}\quad\begin{pmatrix}\lambda^{\frac{1}{2}}&0\\ 0&\lambda^{-\frac{1}{2}}\end{pmatrix},\qquad\text{SCT:}\quad\begin{pmatrix}1&0\\ -c&1\end{pmatrix}$, (31) we get the following commutation relations
@@
Here $j=\frac{1}{2}$, $m=\pm\frac{1}{2}$ and the generators of translation, dilatation and SCT are denoted by $\bar{L}_{-1},\,\bar{L}_{0}$ and $\bar{L}_{1}$ respectively.
The transformation law (30) shows that the doublet $\left(\bar{d}^{\frac{1}{2}}_{\frac{1}{2}}(\bar{z}),\bar{d}^{\frac{1}{2}}_{-\frac{1}{2}}(\bar{z})\right)$ transform in the tensor product of two $SL(2,\mathbb{R})_{\text{R}}$ representations one of which is a conformal highest weight representation with weight $\bar{h}=1/2$ and the other one is a two dimensional spin-$\frac{1}{2}$ representation. Under the spin-$\frac{1}{2}$ representation the fields $\bar{d}^{\frac{1}{2}}_{\pm\frac{1}{2}}(\bar{z})$ have weights $\pm\frac{1}{2}$. To see this at the level of infinitesimal transformation law, we denote the generator in the highest weight ($\bar{h}=\frac{1}{2}$) representation by $\bar{L}_{n}^{\bar{h}}$ and the generator in the spin-$\frac{1}{2}$ representation by $\bar{L}_{n}^{\bar{\sigma}}$. Now one can easily verify that the following commutation relations
@@ VII Correlation functions of dipole currents and matter fields
The dipole operator $\bar{d}^{\frac{1}{2}}_{\alpha}$ has scaling dimension $1/2$. Therefore conformal invariance determines the two point function to be  [^10] In terms of the original field $\bar{d}(\bar{w},\bar{z})$ this is nothing but $\langle\bar{d}(\bar{w}_{1},\bar{z}_{1})\bar{d}(\bar{w}_{2},\bar{z}_{2})\rangle=k\frac{\bar{w}_{1}-\bar{w}_{2}}{\bar{z}_{1}-\bar{z}_{2}}$. This form of the two point function is uniquely determined by the $SL(2,\mathbb{R})_{\text{R}}$ transformation law (28).
@@
where $k$ is a real number which can be zero also. However, as we will see we can take $k$ to be nonzero without changing the single soft photon theorems. Our current understanding of the role of dipole symmetry in the loop level soft theorems does not allow us to determine the value of $k$. This is discussed in more detail in section XI.